# Cislunar Constellation Design for Space Situational Awareness with Time-Expanded p-Median Problem

### v2

@@ 5.3.2 Inter-Orbit Neighborhood
The inter-orbit neighborhood of $j$ is defined as the set of candidate facility locations that are (1) on different LPOs with the same $M$:$N$ resonance, and (2) exhibit a similar solar phase angle $\phi_{S}$ with respect to a reference target at a reference epoch. The latter condition is required to isolate a single slot from the slots along an LPO with the same resonance as the LPO along which slot $j$ is located.
@@ 5.5.3 Memoization
During the heuristic process of the Lagrangian method, it is possible that the Lagrangian relaxed solution $\bar{Y}$ or the result of either an intra- and inter-orbit neighborhood swap result in a choice of facilities that has already been considered by the algorithm. We make use of function memoization on the feasible allocation construction Algorithm 2 by storing the outputs $\underline{X}$ and $\underline{\theta}$ that results from a given set of active facility indices $\mathcal{J}^{\bar{Y}}$.
@@ 6 Numerical Results
The two sets of demand presented in Section 3 are considered. The corresponding problem dimensions are summarized in Table 3. For each set of demand introduced in Section 3, we consider LPOs given in Table 7, with an approximate time-spacing parameter $\Delta t_{b}=12$ hours. Using equation (13), this results in $n=1212$ candidate facility slots in total, summarized in Table 7. Since the set of candidate observer LPOs is assumed to be resonant, any architecture involving any combination of LPOs will return to its original configuration after $N_{\rm syn}$ synodic period, as introduced in Section 2.3.2. Leveraging this feature, we consider a TE-$p$-MP with a time horizon of $N_{\rm syn}=4$ synodic months. Each synodic month is discretized with a time-step of 30 steps per synodic month, corresponding to roughly 24 hours, thus resulting in $\ell=120$ time-steps. Note that the $N_{\rm syn}$-synodic period-repeating feature of the candidate LPOs is not necessitated by the TE-$p$-MP formulation itself; one may choose to formulate the problem with non-repeating candidate observer locations, and set a longer time-horizon as well.
@@ 6.2.1 Cone of Shame
Figure 10 shows the LM-based solutions for $p=2,3,4,5$ and $\bar{m}_{\mathrm{crit}}=20$. In the Cone of Shame demand, the 1:1 resonant L1 Lyapunov orbit is found to be used in many competitive instances. As an exception, when only two observers are placed, an alternative constellation configuration involving also a 3:1 resonant halo orbit is found. In these three solutions, as the number of observers $p$ is increased, an additional observer is added to the same orbit, increasing coverage fraction as well.
@@
The steering history for the selected architectures in Figure 10 are shown in Figure 11. For $p=2$, as shown in Figure 11(a), observer 1 located on the 3:1 resonant Halo steers its sensor both along the azimuth and elevation angles, while observer 2 has its sensor lied on the $xy$-plane. The azimuth steering angle of the two observers exhibit periodicities according to their orbital periods. We note that observer 1 is not disadvantaged by the fact that that it is not on a 1:1 resonance as its sensor is pointed out of plane with a non-zero elevation angle for the majority of time. To provide physical intuition, we plot the steering history of the $p=2$ architecture in position space in Figure 12. For the sake of clarity, we only show the first 10 steps, corresponding to one revolution along observer 1 located on a 3:1 resonant LPO. With architectures for $p=3,4,5$, we are able to see steering histories of observers on the same, 1:1 resonant L1 Lyapunov. The azimuth angle history for each observer follows a similar track shifted according to their location along the LPO. The appropriate azimuth angle to steer the sensor on the 1:1 resonant LPO is largely dictated by the Sun’s illumination direction, causing the observers to follow a similar in-plane steering strategy. Meanwhile, the elevation steering history is distinct for each observer; the out-of-plane sensor pointing direction is varied in order to achieve a more comprehensive coverage of the volume of interest.
@@
Figure 10: Optimized observer locations for Cone of Shame demand at time-step $t=1$ (at a new Moon). Red shaded volume represents region of interest.
@@
Figure 12: Sensor pointing directions of each observer over 10 time-steps for $\bar{m}_{\rm crit}=20$ $p=2$. Red shaded volume represents the region of interest.
@@ 6.2.2 Low-Energy Transfer Transit Window
Figure 14 shows the sensor steering history for architectures in Figure 13. In all cases, all observers are located on one side of the region of interest; as a result, the azimuth angle remains mostly between $-30^{\circ}$ and $30^{\circ}$. With the $p=2$ case utilizing Southern halos, the elevation angle across observers and across time is skewed on the positive side, while with the $p=3,4,5$ cases utilizing Northern halos, the elevation angle is skewed on the negative side. Figure 15 shows the sensor pointing directions for $\bar{m}_{\rm crit}=20$ and $p=4$ from Figure 13(c). As the Northern halo has a larger out-of-plane amplitude than the region of interest, sensors are pointed downward around the apolunes along the Northern halo. We also note that over the first 10 time-steps, the steering directions of each observer have slight variations between one another due to the variation in their respective phase along the LPO.
@@
Figure 13: Optimized observer locations for LET transit demand at time-step $t=1$ (at a new Moon). Red shaded volume represents region of interest.
@@
Figure 15: Sensor pointing directions of each observer over 10 time-steps for $\bar{m}_{\rm crit}=20$ $p=4$. Red shaded volume represents the region of interest.
@@ 6.3 Key Takeaways on the Architecture Design Space
  • For a set of demand centered around L2, LPOs covering a smaller spatial region, in particular 2:1 and 3:1 resonant L2 halos, are favored. These LPOs are suitable as the observer fly in region near but outside the volume of interest; the latter consideration is crucial when considering FOV, as residing inside the volume of interest would severely limit the fraction of volume that is visible to a given observer, despite its proximity to targets.

### v3

@@ 5.3.2 Inter-Orbit Neighborhood
The inter-orbit neighborhood of $j$ is defined as the set of candidate facility locations that are (1) on different LPOs with the same $M$:$N$ resonance, and (2) exhibit a similar solar phase angle $\phi_{S}$ with respect to a reference target at a reference epoch. The inter-orbit neighborhood for the same original slot as in Figure 9(a) is shown in Figure 9(b). The latter condition is required to isolate a single slot from the slots along an LPO with the same resonance as the LPO along which slot $j$ is located.
@@ 5.5.3 Memoization
During the heuristic process of the Lagrangian method, it is possible that the Lagrangian relaxed solution $\bar{Y}$ or the result of either an intra- and inter-orbit neighborhood swap result in a choice of facilities that has already been considered by the algorithm. We make use of function memoization on the feasible allocation construction Algorithm 2 by storing the outputs $\underline{X}$ and $\underline{\theta}$ that result from a given set of active facility indices $\mathcal{J}^{\bar{Y}}$.
@@ 6 Numerical Results
The two sets of demand presented in Section 3 are considered. The corresponding problem dimensions are summarized in Table 3. For each set of demand introduced in Section 3, we consider LPOs given in Appendix 7.1, with an approximate time-spacing parameter $\Delta t_{b}=12$ hours. Using equation (13), this results in $n=1212$ candidate facility slots in total, also presented in Appendix 7.1. Since the set of candidate observer LPOs is assumed to be resonant, any architecture involving any combination of LPOs will return to its original configuration after $N_{\rm syn}$ synodic period, as introduced in Section 2.3.2. Leveraging this feature, we consider a TE-$p$-MP with a time horizon of $N_{\rm syn}=4$ synodic months. Each synodic month is discretized with a time-step of 30 steps per synodic month, corresponding to roughly 24 hours, thus resulting in $\ell=120$ time-steps. Note that the $N_{\rm syn}$-synodic period-repeating feature of the candidate LPOs is not necessitated by the TE-$p$-MP formulation itself; one may choose to formulate the problem with non-repeating candidate observer locations, and set a longer time-horizon as well.
@@ 6.2.1 Cone of Shame
Figure 10 shows the LM-based solutions for $p=2,3,4,5$ and $\bar{m}_{\mathrm{crit}}=20$. In the Cone of Shame demand, the 1:1 resonant L1 Lyapunov orbit is found to be used in many competitive instances. As an exception, when only two observers are placed, an alternative constellation configuration involving also a 3:1 resonant halo orbit is found. In these three solutions, as the number of observers $p$ is increased, an additional observer is added to the same orbit, increasing the coverage fraction as well.
@@
The steering history for the selected architectures in Figure 10 are shown in Figure 11. For $p=2$, as shown in Figure 11(a), observer 1 located on the 3:1 resonant Halo steers its sensor both along the azimuth and elevation angles, while observer 2 has its sensor lied on the $xy$-plane. The azimuth steering angle of the two observers exhibits periodicities according to their orbital periods. We note that observer 1 is not disadvantaged by the fact that it is not on a 1:1 resonance as its sensor is pointed out of plane with a non-zero elevation angle for the majority of the time. To provide physical intuition, we plot the steering history of the $p=2$ architecture in position space in Figure 12. For the sake of clarity, we only show the first 10 steps, corresponding to one revolution along observer 1 located on a 3:1 resonant LPO. With architectures for $p=3,4,5$, we are able to see steering histories of observers on the same, 1:1 resonant L1 Lyapunov. The azimuth angle history for each observer follows a similar track shifted according to their location along the LPO. The appropriate azimuth angle to steer the sensor on the 1:1 resonant LPO is largely dictated by the Sun’s illumination direction, causing the observers to follow a similar in-plane steering strategy. Meanwhile, the elevation steering history is distinct for each observer; the out-of-plane sensor pointing direction is varied in order to achieve a more comprehensive coverage of the volume of interest.
@@
Figure 10: Optimized observer locations for Cone of Shame demand at time-step $t=1$ (at a new Moon)
@@
Figure 12: Sensor pointing directions over 10 time-steps with $\bar{m}_{\rm crit}=20$ $p=2$ for Cone of Shame demand
@@ 6.2.2 Low-Energy Transfer Transit Window
Figure 14 shows the sensor steering history for architectures in Figure 13. In all cases, all observers are located on one side of the region of interest; as a result, the azimuth angle remains mostly between $-30^{\circ}$ and $30^{\circ}$. With the $p=2$ case utilizing Southern halos, the elevation angle across observers and across time is skewed on the positive side, while with the $p=3,4,5$ cases utilizing Northern halos, the elevation angle is skewed on the negative side. Figure 15 shows the sensor pointing directions for $\bar{m}_{\rm crit}=20$ and $p=4$ from Figure 13(c). As the Northern halo has a larger out-of-plane amplitude than the region of interest, sensors are pointed downward around the apolunes along the Northern halo. We also note that over the first 10 time steps, the steering directions of each observer have slight variations from one another due to the variation in their respective phase along the LPO.
@@
Figure 13: Optimized observer locations for LET transit demand at time-step $t=1$ (at a new Moon)
@@
Figure 15: Sensor pointing directions over 10 time-steps with $\bar{m}_{\rm crit}=20$ $p=4$ for LET transit demand
@@ 6.3 Key Takeaways on the Architecture Design Space
  • For a set of demands centered around L2, LPOs covering a smaller spatial region, in particular 2:1 and 3:1 resonant L2 halos, are favored. These LPOs are suitable as the observer fly in the region near but outside the volume of interest; the latter consideration is crucial when considering FOV, as residing inside the volume of interest would severely limit the fraction of volume that is visible to a given observer, despite its proximity to targets.
@@ 7 Conclusions
Acknowledgments
Artificial Intelligence technologies were used for grammar checking.